OpenStudy (anonymous):

Factor the expression completely. x^2-5w+6 how do i do this?

hero (hero):

First find ac = 6 a + c = -5

hero (hero):

Find the value of a and c Have any idea?

hero (hero):

Okay, well a = -2 b = -3 (-2)(-3) = 6 -2 + (-3) = -5

hero (hero):

Now that we know this, we can use that to split the middle term so x^2 - 5w + 6 becomes: x^2 - 2w - 3w + 6 Now we have four factors so we can factor by grouping.

hero (hero):

Oh, wait a minute. I just noticed that you have x and w mixed. You can only have one variable here in order for it to be solveable. Is the original problem w^2 -5w + 6 ?